Calculate how many grams of sodium acetate are needed to make 100.0 ml of acetic acid/sodium acetate buffer with the pH of 5.00 using 0.100 M acetic acid and solid sodium acetate.
Answer #1

PH = 5

PKa = 4.75

no of moles of CH3COOH    = molarity * volume in L

                                          = 0.1*0.1 = 0.01moles

PH   = Pka + log[CH3COONa]/[CH3COOH]

    5      = 4.75 + log[CH3COONa]/0.01

log[CH3COOH]/0.01   = 5-4.75

log[CH3COONa]/0.01    = 0.25

[CH3COONa]/0.01       = 10^0.25

[CH3COONa]/0.01        = 1.7783

[CH3COONa]                 = 1.7783*0.01

[CH3COONa]                 = 0.017783 moles

mass of CH3COOH   = no of moles * gram molar mass

                                   = 0.017783 *82   = 1,46g >>>>answer